OpenStudy (anonymous):

Will medal if right Last one !! In May, Liam and Charlie had the same amount of money in their savings accounts. In June, Liam deposited $140 into his account. Charlie said he increased the money in his account by 5%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in May?

OpenStudy (anonymous):

So it's just asking: "140 is 5% of what?" ie: \[140=0.05x\]\[x=\frac{140}{0.05}\]\[x=$2800\]

OpenStudy (anonymous):

a=b a + 140 b + 0.05b a+140 = a + 0.05a 140 = 0.05a 2800 = a

OpenStudy (anonymous):

im not totally sure this is correct but i think this is the way to solve it amt of money Liam = a Charlie = b a=b a + 140 b + 0.05b a+140 = a + 0.05a 140 = 0.05a 2800 = a

OpenStudy (anonymous):

ty guys everyone give each other medals !!!! @I_Need_Help_With_Home @H3LPN33DED @twvogels

OpenStudy (anonymous):

@H3LPN33DED and @I_Need_Help_With_Home are wright

OpenStudy (anonymous):

Thank you @humphriescody

OpenStudy (anonymous):

Thanks bud @humphriescody